Look at the graph provided below. Consumer surplus is: 18 16 14 12 10 8 6 4 2 St 5 10 15 20 25 30 35 40 45 and has an upward slope. D1 starts at approximately (0,16) and has a downward slope. The two lines cross at the point (30,14)) $60 $120 $30
Answer #1

Equllibrium occurs at (30,14) and the Demand is at (0,16)

Consumer surplus = (1/2) (Change in price) (Change in quantity)

Change in price = (16-14) = 2

Change in quantity = 30-0 = 30

Consumer surplus = (1/2)*2*30

Consumer surplus = 30$